the 1-17. Water has a density of 1.94 slug/ft. What is the density expressed in SI units? Express the answer to three significant figures.
2-3. Determine the magnitude of the resultant force Fe Fi + F2 and its direction, measured counterclockwise from the positive x axis. F, = 250 lb 30 45° F2 375 lb Prob. 2-3
Answer #1

Given water density is1.94 slug/ft^3

Let convert to SI form

= (1.94 slug/ ft^3) × ( ft/0.3048m)^3 × ( 14.59 kg / 1 slug)

= (1.94 × 14.59 )/ 0.3048^3 kg/m^3

=28.3046/ 0.028317 kg/m^3

=999.5675 kg/m^3

=999.5675/1000 Mg/m^3

=0.99956 Mg/m^3

=1.000 Mg/m^3